Two masses, one is m1=2.0 kg and another one m2 = 3.0 kg are connected by a thin string running over a massless pulley. The mass m2 is hanging above the ground and pulling another mass over a 57.27 degree ramp with acceleration 0.7 m/s2. calculate mu_k.
Answer #1

m2g - T = m2a

T - m1gsin(57.27)- um1gcos(57.27) = m1a

=> m2g - m1gsin(57.27)- um1gcos(57.27) = (m1+m2)a

=>  m2g - m1gsin(57.27) - (m1+m2)a = um1gcos(57.27)

=> u = ( m2g - m1gsin(57.27) - (m1+m2)a )/ (m1gcos(57.27))

=> u = (3*9.8 - 2*9.8*0.841 - (2+3)*0.7 ) /( 2*9.8*0.540) = 0.88